(Today's troubleshooting dealt with repair and cleaning of a threaded hole on a standoff element for the right K-BAR capture fitting. The FE-2 used a vacuum cleaner to remove FOD (Foreign Object Debris) plus safety goggles, rubber gloves and a surgical mask for his protection.)
Leo also continued COL commissioning, today unlocking (but not completely removing) the AVM (Anti-Vibration Mount) locking bolts of the module's ISFA (Intermodular Ventilation Supply Fan Assembly) and IRFA (Intermodular Ventilation Return Fan Assembly). (The two fan assemblies are located at opposite sidewalls of the module, both behind cover panels.)
Later, Eyharts assisted FE-1 Malenchenko in modifying (retrofitting) an air duct in the Service Module (SM) with an acoustic shield. (With a joint crewtime of ~3 hrs allotted, work steps included establishing access to the air duct in the SM 'ceiling' by removing a panel (#322), cutting the acoustic shield to size from a larger plate of the soft material with a hacksaw, installing the shield padding on panel 322, re-attaching the panel over the air duct and finally covering the acoustic shield with a piece of protective wire screen.)
In the Lab, CDR Whitson meanwhile performed troubleshooting on ER3 (EXPRESS Rack 3) that earlier (2/25) was found to have a small leak (during demating) at the QD (Quick Disconnect) of the MTL (Moderate Temperature Loop) supply jumper hose between ER3 and the Z-panel. (Troubleshooting steps included fully demating the hose, determining which side of the QD (Z-panel-side or jumper-side) had the problem if still present and, if so, examining the QD and cycling its movable parts if they are stuck. ER3 is slated to be transferred to the Columbus module as part of NASA's share in its utilization (~50%).)
In preparation for 1J/A EVA-5, Whitson had time set aside for taking torque measurements (with a Dial Torque Wrench) on three bolts of the spare TBA-5 (Trundle Bearing Assembly #5). (Accurate running torque values are needed to determine final bolt torque settings for the fifth spacewalk, during which EV1 Bob Behnken and EV2 Mike Foreman will install the TBA in the starboard SARJ (Solar Alpha Rotary Joint).)
After demating the BITS2-12 onboard telemetry measurement system in the SM from the BITS2-12 onboard transmitter (set #1, BP1A), FE-1 Malenchenko removed and replaced the BPA1 with a new unit brought up by Progress 28P. BITS2-12 was remated afterwards. (The first inspection of the transmitter, located between panels 309 & 312, for possible malfunction was conducted by Yuri on 12/6/07.)
The crew had an hour to review transfer and stowage plans for the 1J/A docked period, going through uplinked reference material (1J/A Transfer List, 1J/A Transfer Choreography, 1J/A Stowage Plan, and an overview of Stowage Requirements for the 1J/A Stage of Increment 16.). Afterwards, Peggy, Yuri and Leo held a 30-min. teleconference with ground specialists via S-band/audio to discuss transfer particulars.
In the Lab, Peggy Whitson closed down the OGS (Oxygen Generation System). (This involved deactivating the WDS (Water Delivery System), turning off the OGS in its rack and demating the O2 outlet QD, followed by purging the OGS H2 (Hydrogen) sensor and re-installing the WGS on the OGS rack front.)
With both OGS and the Elektron-VM O2 (Oxygen) generator currently off, cabin air refreshes with O2 are being performed by Malenchenko from Progress 28P storage (SrPK) as required.
Using the Russian IPD Draeger tube sampler (sensors 5C & 2A), Malenchenko checked for CO (Carbon Monoxide) in the SM.
Also in the Zvezda module, the FE-1 afterwards took readings of potentially harmful atmospheric contaminants with the CMS (Countermeasure System), part of the GANK-4M Real-Time Harmful Contaminant Gas Analyzer suite, which uses preprogrammed microchips to measure H2CO (Formaldehyde, methanal), CO and NH3 (Ammonia), taking one measurement per microchip;
Continuing preparations in the U.S. Airlock for the five 1J/A EVAs, Whitson first inspected APFR WIF (Articulated Portable Foot Restraint/Worksite Interface Fixture) adapters for sharp edges, to aid in the ongoing spacesuit glove cut incident investigation, then terminated the BSA (Battery Stowage Assembly) recharging of the first set of EMU (Extravehicular Mobility Unit) batteries and started recharge on the second set.
As part of the current round of periodic preventive maintenance of RS (Russian Segment) ventilation systems, Yuri spent 20 min. in the Soyuz TMA-11/15S at the FGB nadir port, cleaning the screen of its BVN fan/heater assembly.
Afterwards, the Russian flight engineer conducted the periodic (monthly) functional closure test of the Vozdukh CO2 removal system's spare emergency vacuum valves (AVK), in the spare parts kit. (The AVKs are critical because they close the Vozdukh's vacuum access lines in the event of a malfunction in the regular vacuum valves (BVK) or a depressurization in the Vozdukh valve panel (BOA). Access to vacuum is required to vent CO2 during the regeneration of the absorbent cartridges (PP). During nominal operation, the AVK valves remain open.)
To support ground-conducted troubleshooting of the TVIS treadmill, CDR Whitson today had a TVIS checkout activity added to her timeline. After undergoing its regular 6-month maintenance on 3/3, a TVIS Stabilization Fault (S3) appeared during its first use, rendering the exercise device off limits for the crew all day yesterday. (If there was no recurrence of the stabilization fault during today's checkout, Peggy was to perform a speed characterization test and download the test data as part of the nominal exercise file. The crew was then Go for using the TVIS as planned. If not, the treadmill would remain off limits until further troubleshooting, and the crew would use the other exercise hardware.)
The FE-1 completed the routine maintenance of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including ASU toilet facilities systems/replaceables. (Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, replacement of the KOV EDV for the Elektron-intended water, and processing U.S. condensate water as it becomes available in a filled CWC (Contingency Water Container) from the Lab humidifier.)
Yuri also conducted the daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur).
In addition, working off his voluntary 'time permitting' task list, Yuri -
Conducted the regular daily checkup on the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at +20 degC, including a temperature check on its ART (automatic temperature recorder), and
Performed the daily monitoring, picture-taking and downloading for the BIO-5 Rasteniya-2 ("Plants-2") experiment which researches growth and development of plants (peas) under spaceflight conditions in the Lada-12 greenhouse from IBMP (Institute of Bio-Medical Problems, Russian: IMBP).
At ~4:45am EST, Leo Eyharts powered up the SM's amateur radio equipment (Kenwood VHF transceiver with manual frequency selection, headset, and power supply), to conduct, at 4:50am, a ham radio exchange with students at the LycĂ(c)e (Secondary School) de la Borde Basse in Castres, France. Questions to Leo were uplinked by ARISS (Amateur Radio on ISS) beforehand. ('What type of schooling and higher studies did you complete to enable you to become an astronaut? How long did you study for it?'; 'What is the typical duration for a mission on board the space station (minimum & maximum?)'; 'Could you describe one of the experiments that you are to conduct in the Columbus laboratory?'; 'How do you resolve cases of onboard personal conflict (arguments)?'; 'Do you lose the notion of time on the ISS? Do you have the same sleeping patterns as when you are on Earth?')
At ~7:35am, the ISS crew convened for their weekly teleconference with ISS Program Management at JSC/Houston via Private S/G2, S-band/audio.
COL Anomalies: As reported here yesterday, the FSL (Fluid Science Laboratory) rack shut down shortly after start-up and remained powered-down. Troubleshooting by Col-CC (Columbus Control Center/Oberpfaffenhofen) was performed today after Eyharts installed a NASA videocamera in front of FSL and checked a UIP (Utility Interface Panel) connector. The EuTEF (European Technology Exposure Facility) payload's Earth-viewing camera has failed to start up twice and is currently off. The SOLAR (Solar Monitoring Observatory) payload had a spontaneous reboot Monday (3/3) night and came back to nominal performance afterwards; Col-CC is assessing.
BCC Checkout Failure: Yesterday (3/3-3/4 overnight), ground teams initiated the standard two-hour checkout of BCC (Backup Control Center) swing and activation procedures that would be necessary in the event of a flight control transfer from Houston to the HSG (Houston Support Group) at TsUP-Moscow. The checkout failed when RGS-34 (Russian Groundsite 34) was unable to uplink a PPCP (PrePlanned Command Package) from MCC-Houston on DO3. For lack of time, the backup pass uplink on DO5 was cancelled as well. HGS-M plans to re-perform the BCC Checkout at end-March. (Purpose of the BCC Checkout is to demonstrate BCC functionality and provide proficiency training for HSR (Houston Support Room) personnel.)
CEO photo targets uplinked for today were Lahore, Pakistan (ISS had a nadir pass at midday over the Pakistan's second largest city of over 7 million people, approaching from the NW. As the station broke out of the mountains of Afghanistan, the crew was to look for this target on the northeastern part of the Indus River Plain, using the long lens settings and trying to map the urban margins), Mount Vesuvius (this famous stratovolcano is located on the west coast of Italy just east of the city of Naples. ISS pass was late morning southeastward down the Italian peninsula, and the crew was to aim just right of track for the bay of Naples and this isolated peak, using the long lens settings for details of the structure), and Madrean Sky Islands (this target is located in the northern reaches of Mexico's Sierra Madre Occidental which boast some of the richest biodiversity anywhere in North America. It is a veritable archipelago of cool, moist, higher-altitude pine-oak forested mountain ranges that dot the hot, lower Sonoran and Chihuahuan deserts of southern Arizona and New Mexico and northwestern Mexico. These climatological islands are situated in remote and rugged areas and are vestiges of cooler, wetter periods during the ice ages. On this midday pass from the NW, the crew was to use the short lens for a nadir, contextual mapping strip across the heart of this region).